OpenStudy (anonymous):

A culture started with 6,000 bacteria. After 3 hours, it grew to 7,200 bacteria. Predict how many bacteria will be present after 19 hours. Round your answer to the nearest whole number.

OpenStudy (amistre64):

i believe bacteria follow the continuous model Pe^(rt)

OpenStudy (amistre64):

the rate here is what important, or rather what is missing A = Pe^rt A/P = e^rt ln(A/P) = rt ln(A/P)/t = r would be the general formula; and just fill in the specifics

OpenStudy (anonymous):

Idk how :(((

OpenStudy (amistre64):

what part of it cant you fill in?

OpenStudy (anonymous):

I am very unsuccessful when it comes to forming an equation out of a word problem in general.

OpenStudy (amistre64):

i used the convention for finances; perhaps we could rename that parts for you....

OpenStudy (anonymous):

All I know that this is exponential growth, and I have no clue how do do anything.

OpenStudy (amistre64):

P = starting amount; lets call it "Sa" A = ending amount; lets call it "Ea" rate = rate, and time = time so those can stay the same name ln(Ea/Sa)/t = r would be a decent translation

OpenStudy (amistre64):

What are our stated values then for Sa, Ea, r, and t?

OpenStudy (anonymous):

6000*(1.2)^(19/3) = 19039

OpenStudy (anonymous):

A culture started with 2,000 bacteria. After 4 hours, it grew to 2,200 bacteria. Predict how many bacteria will be present after 9 hours.

OpenStudy (amistre64):

thats not what you have posted above: "A culture started with 6,000 bacteria. After 3 hours, it grew to 7,200 bacteria. Predict how many bacteria will be present after 19 hours"

OpenStudy (anonymous):

I know, the next question was the same problem with different numbers. :P

OpenStudy (anonymous):

2000*(1.1)^(9/4)= 2478

OpenStudy (amistre64):

that is why the process is more important to know that any answers to it

OpenStudy (amistre64):

the process never changes; just the numbers you apply to it

OpenStudy (anonymous):

amistre can you help me with exponential decay? PLEASE <3

OpenStudy (anonymous):

Sometimes there are things going on that others on the outside can't understand, I'm sorry amistre64, you've helped me a great deal with a lot of things, but there are new personal things going on that require me to finish as quickly as possible. I'm not trying to be rude, sorry if I am, but there really is a reason for why answers would help me now than the equations. Sounds ludicrous, I know and understand, but please trust me (even though I'm a stranger).

OpenStudy (anonymous):

I've worked very hard to get where I am though.

OpenStudy (amistre64):

ill trust you :) and i wish you all the luck

OpenStudy (anonymous):

Thank you.